OpenStudy (anonymous):

The sum of three consecutive odd integers is 177. What are the integers?

hartnn (hartnn):

let the smallest odd integer be x. so next 2 integers will be x+2 and x+4. and since their sum =177 x+x+2+x+4=177 can u solve this to find x ?

OpenStudy (anonymous):

if its x+x+2+x+4=177 what am i looking for with the x? how do i find x if theres several and how do i find x if there isnt an x on the 177's side?

hartnn (hartnn):

u just combine like terms so here x+x+x = 3x ok? 2+4 = 6 this was left side, 3x+6 right side 177. so u have 3x+6 = 177 now can u solve further ?

hartnn (hartnn):

@qnqwiz what did u get the value of x as ?

OpenStudy (anonymous):

let me try 1 sec

hartnn (hartnn):

okay :)

OpenStudy (anonymous):

let the three integers be x,x+2 and x+4 hence ,3x +6=177 3x =177-6=171 x=171/3=57 so ,the integers are 57,57+2=59 and 57+4=61

OpenStudy (anonymous):

i got 59, 61 and 63, but i cheked my math and had it wrong.got it now though! i added too much

hartnn (hartnn):

but did u finally get 57,59,61 ?

OpenStudy (anonymous):

57+59+61=177, but thats it?

OpenStudy (anonymous):

yes i did

hartnn (hartnn):

thats good :)

OpenStudy (anonymous):

thanks!!

hartnn (hartnn):

welcome :)